# Semi-Supervised Clustering of Sparse Graphs: Crossing the Information-Theoretic Threshold

### 08 Jan 2024 v3

@@ 4.3 Semi-Supervised Detectability
Let $(A,x)$ obey the planted bisection model $G(n,\frac{a}{n},\frac{b}{n})$ and denote $\langle xx^{\top},\tilde{A}\rangle$ as $Y$. Then for any $\epsilon>0$, we have $Y/n\in[\frac{a-b}{2}-\epsilon,\frac{a-b}{2}+\epsilon]$ with probability converging to one as $n\to\infty$.
@@
$\operatorname{Var}Y=4\left(a\left(\frac{n}{4}-\frac{1}{2}\right)\left(1-\frac{a}{n}\right)+b~{}\frac{n}{4}\left(1-\frac{b}{n}\right)\right)\leq n(a+b)$ (95)
Then Chebyshev’s inequality implies that for any $\delta\ in(0,1)$
@@
$\implies\quad\operatorname{P}\left(|\frac{Y}{n}-\frac{a-b}{2}+\frac{a}{n}|\geq\frac{\sqrt{a+b}}{n^{\delta/2}}\right)\leq\frac{1}{n^{1-\delta}}$ (97)
@@
$\operatorname{P}\left(\frac{Y}{n}\geq\frac{a-b}{2}+\epsilon\right)+\operatorname{P}\left(\frac{Y}{n}\leq\frac{a-b}{2}-\epsilon\right)\leq\frac{1}{n^{1-\delta}}$ (99)
@@
$\operatorname{P}\left(\frac{Y}{n}\in[\frac{a-b}{2}-\epsilon,\frac{a-b}{2}+\epsilon]\right)\geq 1-\frac{1}{n^{1-\delta}}$ (100)
@@
$\text{CSDP}(\tilde{A})\geq\langle xx^{\top},\tilde{A}\rangle$ (102)
@@
$T(G,x_{\mathcal{R}};\Delta)=\begin{cases}1\quad\quad&\text{if~{}}\text{CSDP}(\tilde{A})\geq n[(a-b)/2-\Delta]\\ 0\quad\quad&\text{otherwise}\end{cases}$ (103)
The following lemma bounds the CSDP of ERM from above. Intuitively, the contribution of blocks of adjacency matrix, where columns or rows are associated with revealed nodes, concentrates well around zero. So the ‘effective dimension’ of the SDP is reduced, hence the optimal value.
@@
This result is rigorously derived with profound insights from mathematical physics. However, there is an implicit condition on the average degree $d$ in the proof. In fact, it is common to assume at least $d>1$ in the literature concerning unsupervised clustering because otherwise the graph has no giant component, not to mention reconstruction, as discussed in Section 1.2. However, our approach leads to a subgraph with possibly small effective average degree. Moreover, we do not want to be limited by the topology structure, although it is indeed a fundamental limit in the unsupervised setting. Theorem 12 shows that semi-supervised SDPs are able to integrate those sublinear components. To achieve that we resort to Grothendieck’s inequality and carry out the analysis without assumption on $d$.
@@
Here $K_{\text{G}}$ is an absolute constant called Grothendieck’s constant. We consider the upper bound derived in (Braverman et al., 2011),
$K_{\text{G}}<\frac{\pi}{2\log(1+\sqrt{2})}\leq 1.78$ (107)
Notice that if we restrict the vectors $X_{i}$’s and $Y_{i}$’s to the unit sphere $\mathcal{S}^{n-1}$, the inequality still holds. Since $s,t$ are arbitrary, the left hand side of inequality (105) is indeed the $\ell_{\infty}\to\ell_{1}$ norm of matrix $M$, which is
$\|M\|_{\infty\to 1}=\max_{\|x\|_{\infty}\leq 1}\|Mx\|_{1}=\max_{s,t\in\{-1,1\}^{n}}s^{\top}Mt=\max_{s,t\in\{-1,1\}^{n}}\big{|}\sum_{i,j}M_{ij}s_{i}t_{j}\big{|}$ (108)
This norm is also known as the cut norm, whose importance in algorithmic problems is well understood in theoretical computer science community. With the elliptope definition of SDP from equation (68) and the consequential factorization of $X$ in equation (71), we can rewrite the theorem in the following matrix form.
@@
$\operatorname{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}>t\right)\leq\exp\left(-\frac{nt^{2}/2}{\sigma^{2}+\frac{Mt}{3}}\right)$ (110)
@@
$\operatorname{P}\left(\frac{2}{n(n-1)}\sum_{i<j}X_{ij}>t\right)\leq\exp\left(-\frac{n(n-1)t^{2}/4}{\frac{4d}{n}+\frac{2t}{3}}\right)$ (115)
@@
$\operatorname{P}\left(\sum_{i<j}X_{ij}>6(1+d)n\right)\leq\exp\left(-3(n-1)\right)$ (116)
@@
$\operatorname{P}\left(\max_{s,t\in\{-1,1\}^{n}}\sum_{i<j}(A-\operatorname{E}A)_{ij}~{}(s_{i}t_{j}+s_{j}t_{i})>6(1+d)n\right)\leq 2^{2n}\cdot e^{-3(n-1)}$ (117)
We conclude the proof with the identity of $\ell_{\infty}\to\ell_{1}$ norm and the fact that theright hand side of the above inequality is less than $5^{-n+2}$.
@@
$\bar{X}^{(n)}=\frac{X_{1}^{(n)}+X_{2}^{(n)}+\dots+X_{n}^{(n)}}{n}$ (118)
then for any $\epsilon>0$,
@@
$\leq\sigma^{2}/n$ (by uniform boundedness) (122)
@@
$\implies$ $\operatorname{P}\big{(}\bar{X}^{(n)}\geq\mu+\epsilon\big{)}\leq\frac{\sigma^{2}}{n\epsilon^{2}}$ (125)
@@
Comparing with a standard large deviation theory, this result allows the random variables to not be identically distributed. And more importantly, the distributions can depend on $n$. Furthermore, the random variables associated with different $n$ are not necessary to be independent.
@@
Proof  Notice that for any feasible $X$ of above optimization problem, we have $X\succeq 0,X_{ii}=1~{}\forall i\in[n+1]$. So, for any $i,~{}j\in[n+1]$,
$(\mathbf{e}_{i}\pm\mathbf{e}_{j})^{\top}X(\mathbf{e}_{i}\pm\mathbf{e}_{j})=2\pm 2X_{ij}\geq 0\quad\implies\quad|X_{ij}|\leq 1$ (128)
@@
$\operatorname{E}|Z|\leq\operatorname{E}(Z^{2})=\operatorname{Var}Z\leq d\frac{m}{n}$ (132)
$\operatorname{Var}|Z|=E(Z^{2})-(\operatorname{E}|Z|)^{2}\leq\operatorname{Var}Z\leq d\frac{m}{n}$ (133)
@@
$\bar{X}^{(n)}\coloneqq\frac{\sum_{j\in[n]\setminus\mathcal{R}}|B_{0j}|}{n-m}$ (134)
@@
$\lim_{n\to\infty}\operatorname{P}\left(\frac{1}{n-m}\sum_{j\in[n]\setminus\mathcal{R}}|B_{0j}|>d\frac{m}{n}+\epsilon\right)=0$ (135)
@@
$M^{(\mathcal{R})}_{ij}=\begin{cases}M^{\text{agg}}_{ij}&\quad i=0\text{~{}or~{}}j=0\\ 0&\quad\text{otherwise}\end{cases}$ (138)
@@
$\leq\text{SDP}(\tilde{A}^{(\mathcal{R}^{\complement})})+\text{SDP}(\tilde{A}^{(\mathcal{R})})$ (143)
@@
$=\text{SDP}\left(A_{\mathcal{R}^{\complement}}-\operatorname{E}A_{\mathcal{R}^{\complement}}\right)-(1-\rho)d$ (145)
Now we apply the Grothendieck’s inequality and Lemma 29. With probability at least $1-5^{-(1-\rho)n+2}$,
$\text{SDP}\left(\tilde{A}^{(\mathcal{R}^{\complement})}\right)\leq 6K_{\text{G}}[1+(1-\rho)d](n-m)<12(1+d)(1-\rho)n$ (146)
@@
$\operatorname{P}_{0}(T(G,x_{\mathcal{R}})=1)\to 0\qquad(n\to\infty)$ (148)
@@ 5 Numerical Experiments
Each point in Figure 4 represents one realization of a SBM with $n=1000$. The dashed line stands for the KS and information-theoretic threshold. The graphs are shared by both the unsupervised and the semi-supervised SDPs. Labels are identified by applying the k-means method to corresponding optimizers. Overlaps of unsupervised SDP essentially drop down to zero on the left-hand side. While, with $20\%$ of the labels revealed, the outcome of our constraint SDP algorithm goes down gradually as the SNR decreases and remains substantially greater than zero even when $\text{SNR}\leq 1$.
Figure 5: Overlap heatmaps of the unsupervised (left) and the semi-supervised (right) SDPs. The coordinates correspond to the model parameters $a$ and $b$. The solid line represents the KS and information-theoretic threshold. The dash line corresponds to $a=b$.
The phase transition theory 3 guarantees that the upper left corner of the left image will be totally dark as $n\to\infty$. But we see semi-supervised SDPs successfully ‘light up’ the area between the two reference lines, see Figure 5. Moreover, when $n$ is sufficiently large, there will be no pixel with value 0.
Figure 6 shows color-coded entry values of optimizer $X^{*}$ in different settings and suggests that representing of the underlying community structure is significantly enhanced by the semi-supervised approach, while no such structure is introduced if there should not be one.
Figure 6: Visualization of the optimizer $X^{*}$. The upper row is concerned with one realization of the SBM $\operatorname{\mathcal{G}}(1000,12/1000,5/1000)$, where the left image shows the value of optimizer for the unsupervised SDP and the right image is associated with the semi-supervised SDP with $\rho=0.2$. The lower left image is optimizer for one realization of the ERM of the same size with the associated average degree $d=8.5$, indices of which are reordered such that the entries related to revealed labels are gathered in four corners. It could be understood as the situation of null hypothesis we defined in Section 4.2.
To see how such a better representation leads to a successful test that is originally impossible, we consider the following simulations. We generate $50$ independent realizations of underlying random graphs ($n=200$) and compute their SDP values with and without the semi-supervised constraints ($\rho=0.25$). Particularly, the parameters in Figure 7 are chosen to have $\text{SNR}>1$. The left two boxes imply that we can tell the difference between SBM and the ERM with the same average degree $d=(a+b)/2$. However, as in Figure 8, the vanilla SDP gives essentially the same result since the two models become contiguous if $\text{SNR}\leq 1$. As we have proved in Theorem 12, our semi-supervised SDP algorithm still manages to distinguish them by bringing down the optimal value of ERM more significantly comparing to its effect on SBM, which is confirmed by the right two boxes.
@@ 6 Conclusion
The census method comes from the combinatorial perspective, while the CSDP is inspired by convex optimization research. Both algorithms are computationally efficient. The former has no requirement on the reveal ratio. The latter one is more practical and backward compatible to the unsupervised setting. By carefully integrating the revealed information with the observed graph structure, we can not only improve the performance of clustering algorithms but resolve initially unsolvable problems. The fundamental changes brought by semi-supervised approach let us cross KS threshold, information-theoretical threshold and even the topological limitation.
Our work provides a different angle to study stochastic models of network and semidefinite programs. In real-world situations, it is almost always the case that we will have a certain fraction of samples being understood fairly well. So an abstract model should be able to capture the existence of such knowledge instead of being blindly restricted to unsupervised setting. Combining the universality of ‘revealed’ information and the insight derived from our census method, it is arguable that the phase transitions, although very mathematically beautiful, will never be an issue in practice. Our results on CSDPs, in turn, could be used to study SDPs, e.g. prove or disprove it can reach the phase transition threshold or the monotone-robustness threshold by a limiting process of $\rho\to 0$.

### 28 Feb 2024 v4

@@ 4.3 Semi-Supervised Detectability
Let $(A,x)$ obey the planted bisection model $G(n,\frac{a}{n},\frac{b}{n})$ and denote $\langle xx^{\top},\tilde{A}\rangle$ as $Y$. Then, for any $\epsilon>0$, we have $Y/n\in[\frac{a-b}{2}-\epsilon,\frac{a-b}{2}+\epsilon]$ with probability converging to one as $n\to\infty$.
@@
$\operatorname{Var}Y=4\left(a\left(\frac{n}{4}-\frac{1}{2}\right)\left(1-\frac{a}{n}\right)+b~{}\frac{n}{4}\left(1-\frac{b}{n}\right)\right)\leq n(a+b).$ (95)
Then Chebyshev’s inequality implies that for any $\delta\in(0,1)$,
@@
$\implies\quad\operatorname{P}\left(|\frac{Y}{n}-\frac{a-b}{2}+\frac{a}{n}|\geq\frac{\sqrt{a+b}}{n^{\delta/2}}\right)\leq\frac{1}{n^{1-\delta}}.$ (97)
@@
$\operatorname{P}\left(\frac{Y}{n}\geq\frac{a-b}{2}+\epsilon\right)+\operatorname{P}\left(\frac{Y}{n}\leq\frac{a-b}{2}-\epsilon\right)\leq\frac{1}{n^{1-\delta}}.$ (99)
@@
$\operatorname{P}\left(\frac{Y}{n}\in[\frac{a-b}{2}-\epsilon,\frac{a-b}{2}+\epsilon]\right)\geq 1-\frac{1}{n^{1-\delta}}.$ (100)
@@
$\text{CSDP}(\tilde{A})\geq\langle xx^{\top},\tilde{A}\rangle.$ (102)
@@
$T(G,x_{\mathcal{R}};\Delta)=\begin{cases}1\quad\quad&\text{if~{}}\text{CSDP}(\tilde{A})\geq n[(a-b)/2-\Delta],\\ 0\quad\quad&\text{otherwise}.\end{cases}$ (103)
The following lemma bounds the CSDP of ERM from above. Intuitively, the contribution of blocks of the adjacency matrix, where columns or rows are associated with revealed nodes, concentrates well around zero. So the ‘effective dimension’ of the SDP is reduced, hence the optimal value.
@@
This result is rigorously derived with profound insights from mathematical physics. However, there is an implicit condition on the average degree $d$ in the proof. It is common to assume at least $d>1$ in the literature concerning unsupervised clustering because otherwise, the graph has no giant component, not to mention reconstruction, as discussed in Section 1.2. But our approach leads to the subgraph with a possibly small effective average degree. Moreover, we do not want to be limited by the topology structure, although it is indeed a fundamental limit in the unsupervised setting. Theorem 12 shows that semi-supervised SDPs are capable of integrating those sublinear components. To achieve that we resort to Grothendieck’s inequality and carry out the analysis without assumption on $d$.
@@
$K_{\text{G}}$ is an absolute constant called Grothendieck’s constant. The inequality was initially proved as a fundamental tool in Functional Analysis. In this paper, we focus exclusively on the above matrix version in $\mathbb{R}^{n}$ and consider the following suboptimal estimate derived by (Braverman et al., 2011),
$K_{\text{G}}<\frac{\pi}{2\log(1+\sqrt{2})}\leq 1.78.$ (107)
Note that if we restrict the vectors $X_{i}$’s and $Y_{i}$’s to the unit sphere $\mathcal{S}^{n-1}$, the inequality is still true as any inequality valid for a set also holds for its subset. Since $s$ and $t$ are arbitrary, the left-hand side of inequality (105) is indeed the $\ell_{\infty}\to\ell_{1}$ norm of matrix $M$, which is
$\|M\|_{\infty\to 1}=\max_{\|x\|_{\infty}\leq 1}\|Mx\|_{1}=\max_{s,t\in\{-1,1\}^{n}}s^{\top}Mt=\max_{s,t\in\{-1,1\}^{n}}\big{|}\sum_{i,j}M_{ij}s_{i}t_{j}\big{|}.$ (108)
This norm is also known as the cut norm, whose importance in algorithmic problems is well understood in the theoretical computer science community. With the elliptope definition of SDP from equation (68) and the consequential factorization of $X$ in equation (71), we can rewrite the theorem in the following matrix form.
@@
$\operatorname{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}>t\right)\leq\exp\left(-\frac{nt^{2}/2}{\sigma^{2}+\frac{Mt}{3}}\right).$ (110)
@@
$\operatorname{P}\left(\frac{2}{n(n-1)}\sum_{i<j}X_{ij}>t\right)\leq\exp\left(-\frac{n(n-1)t^{2}/4}{\frac{4d}{n}+\frac{2t}{3}}\right).$ (115)
@@
$\operatorname{P}\left(\sum_{i<j}X_{ij}>6(1+d)n\right)\leq\exp\left(-3(n-1)\right).$ (116)
@@
$\operatorname{P}\left(\max_{s,t\in\{-1,1\}^{n}}\sum_{i<j}(A-\operatorname{E}A)_{ij}~{}(s_{i}t_{j}+s_{j}t_{i})>6(1+d)n\right)\leq 2^{2n}\cdot e^{-3(n-1)}.$ (117)
We conclude the proof with the identity of $\ell_{\infty}\to\ell_{1}$ norm and the fact that the right-hand side of the above inequality is less than $5^{-n+2}$.
@@
$\bar{X}^{(n)}=\frac{X_{1}^{(n)}+X_{2}^{(n)}+\dots+X_{n}^{(n)}}{n},$ (118)
then, for any $\epsilon>0$,
@@
$\leq\sigma^{2}/n.$ (by uniform boundedness) (122)
@@
$\implies$ $\operatorname{P}\big{(}\bar{X}^{(n)}\geq\mu+\epsilon\big{)}\leq\frac{\sigma^{2}}{n\epsilon^{2}}.$ (125)
@@
Compared with a standard large deviation theory, this result allows the random variables to not be identically distributed. And more importantly, the distributions can depend on $n$. Furthermore, the random variables associated with different $n$ are not necessary to be independent.
@@
Proof  Notice that for any feasible $X$ of the above optimization problem, we have $X\succeq 0,X_{ii}=1~{}\forall i\in[n+1]$. So, for any $i,~{}j\in[n+1]$,
$(\mathbf{e}_{i}\pm\mathbf{e}_{j})^{\top}X(\mathbf{e}_{i}\pm\mathbf{e}_{j})=2\pm 2X_{ij}\geq 0\quad\implies\quad|X_{ij}|\leq 1.$ (128)
@@
$\operatorname{E}|Z|\leq\operatorname{E}(Z^{2})=\operatorname{Var}Z\leq d\frac{m}{n},$ (132)
$\operatorname{Var}|Z|=E(Z^{2})-(\operatorname{E}|Z|)^{2}\leq\operatorname{Var}Z\leq d\frac{m}{n}.$ (133)
@@
$\bar{X}^{(n)}\coloneqq\frac{\sum_{j\in[n]\setminus\mathcal{R}}|B_{0j}|}{n-m}.$ (134)
@@
$\lim_{n\to\infty}\operatorname{P}\left(\frac{1}{n-m}\sum_{j\in[n]\setminus\mathcal{R}}|B_{0j}|>d\frac{m}{n}+\epsilon\right)=0.$ (135)
@@
$M^{(\mathcal{R})}_{ij}=\begin{cases}M^{\text{agg}}_{ij}&\quad i=0\text{~{}or~{}}j=0,\\ 0&\quad\text{otherwise}.\end{cases}$ (138)
@@
$\leq\text{SDP}(\tilde{A}^{(\mathcal{R}^{\complement})})+\text{SDP}(\tilde{A}^{(\mathcal{R})}).$ (143)
@@
$=\text{SDP}\left(A_{\mathcal{R}^{\complement}}-\operatorname{E}A_{\mathcal{R}^{\complement}}\right)-(1-\rho)d.$ (145)
Now we apply Grothendieck’s inequality and Lemma 29. With probability at least $1-5^{-(1-\rho)n+2}$,
$\text{SDP}\left(\tilde{A}^{(\mathcal{R}^{\complement})}\right)\leq 6K_{\text{G}}[1+(1-\rho)d](n-m)<12(1+d)(1-\rho)n.$ (146)
@@
$\operatorname{P}_{0}(T(G,x_{\mathcal{R}})=1)\to 0\quad(n\to\infty).$ (148)
@@
Remark 34
We believe that it is possible to reduce the requirement on $\rho$ through more involved analysis. Empirically, it does not need to be close to 1 for a clear improvement in the result. For instance, our simulation in the next section shows that phase transition disappears with $20\%$ of the nodes revealed.
@@ 5 Numerical Experiments
Each point in Figure 4 represents one realization of a SBM with $n=1000$. The dashed line stands for the KS and information-theoretic threshold. The graphs are shared by both the unsupervised and the semi-supervised SDPs. Labels are identified by applying the k-means method to corresponding optimizers. Overlaps of unsupervised SDP essentially drop down to zero on the left-hand side. In contrast, with $20\%$ of the labels revealed, the outcome of our constraint SDP algorithm goes down gradually as the SNR decreases and remains substantially greater than zero even when $\text{SNR}\leq 1$.
Figure 5: Overlap heatmaps of the unsupervised (left) and the semi-supervised (right) SDPs. The coordinates correspond to the model parameters $a$ and $b$. The solid line represents the KS and information-theoretic threshold. The dashed line corresponds to $a=b$.
Theorem 3 guarantees that the upper left corner of the left image will be completely dark as $n\to\infty$. But we see semi-supervised SDPs successfully ‘light up’ the area between the two reference lines, see Figure 5. Moreover, when $n$ is sufficiently large, there will be no pixel with a value of 0.
Figure 6 shows color-coded entry values of optimizer $X^{*}$ in different settings and suggests that the representing of underlying community structure is significantly enhanced by the semi-supervised approach, while no such structure is introduced if there should not be one.
Figure 6: Visualization of the optimizer $X^{*}$. The upper row is concerned with one realization of the SBM $\operatorname{\mathcal{G}}(1000,12/1000,5/1000)$, where the left image shows the value of optimizer for the unsupervised SDP and the right image is associated with the semi-supervised SDP with $\rho=0.2$. The lower left image is the optimizer for one realization of the ERM of the same size with the associated average degree $d=8.5$, indices of which are reordered such that the entries related to revealed labels are gathered in four corners. It could be understood as the situation of the null hypothesis we defined in Section 4.2.
To see how such a better representation leads to a successful test that is originally impossible, we consider the following simulations. We generate $50$ independent realizations of underlying random graphs ($n=200$) and compute their SDP values with and without the semi-supervised constraints ($\rho=0.25$). Particularly, the parameters in Figure 7 are chosen to have $\text{SNR}>1$. The left two boxes imply that we can tell the difference between SBM and the ERM with the same average degree $d=(a+b)/2$. However, as in Figure 8, the vanilla SDP gives essentially the same result since the two models become contiguous if $\text{SNR}\leq 1$. As we have proved in Theorem 12, our semi-supervised SDP algorithm still manages to distinguish them by bringing down the optimal value of ERM more significantly when compared to its effect on SBM, which is confirmed by the right two boxes.
@@ 6 Conclusion
The census method comes from the combinatorial perspective, while the CSDP is inspired by convex optimization research. Both algorithms are computationally efficient. The former has no requirement on the reveal ratio. The latter one is more practical and backward compatible with the unsupervised setting. By carefully integrating the revealed information with the observed graph structure, we can not only improve the performance of clustering algorithms but also resolve initially unsolvable problems. The fundamental changes brought by the semi-supervised approach let us cross the KS threshold, information-theoretical threshold, and even the topological limitation.
Our work provides a different angle to study stochastic models of networks and semidefinite programs. In real-world situations, it is almost always the case that we will have a certain fraction of samples being understood fairly well. So, an abstract model should be able to capture the existence of such knowledge instead of being blindly restricted to the unsupervised setting. Combining the universality of ‘revealed’ information and the insight derived from our census method, it is arguable that the phase transitions, although mathematically beautiful, will never be an issue in practice. Our results on CSDPs, in turn, could be used to study SDPs, e.g., prove or disprove that it can reach the phase transition threshold or the monotone-robustness threshold by a limiting process of $\rho\to 0$.